OpenStudy (anonymous):

the ratio of boys to girls in the auditorium was 4 to 5 if there were 40 boys in the auditorium how many girls were there

OpenStudy (anonymous):

50

OpenStudy (anonymous):

b=boys g=girls b/g = 4/5 b=40, so: 40/g = 4/5 so, 4g=5*40 g=50 50 girls

OpenStudy (anonymous):

for 4 boys there are 5 girls :P so for 40 boys there are 50 girls